# Mapping global offshore wind resource: wake losses, optimisation potential and climate effects

### v1

@@ 4 Conclusions
In this study, we have investigated global spatial variation in (i) ‘ambient’ offshore wind power potential, and (ii) intra-farm wake losses for a ‘standard’ offshore wind farm with a design representative of near-future installations. The spatial distribution of ambient wind power potential follows well-established global wind climatology, with resources greatest in the mid-latitudes, and lowest in the tropics. We find a strong correlation between ambient resources and percentage wake-induced losses; regions with greater ambient resource tend to lose a smaller fraction of their ambient power to wake effects. This trend exacerbates the spatial variation in ambient wind power density.
@@
However, there remains a broad distribution of wake-induced losses even amongst locations with similar ambient resources. We find that the tropics are particularly badly impacted by wake effects, with the mid-latitudes relatively less impacted. That is, for a tropical and mid-latitude location with the same ambient wind resource, the mid-latitude location is likely to suffer from smaller wake-induced losses than the tropical location. We find that this can be explained by the specific wind speed distributions found in each region. The tropics tend to exhibit narrower wind speed distributions than mid-latitude regions, and thus a greater proportion of the wind speed distribution falls between the cut-in and rated speeds of the wind turbine, where power outputs are most sensitive to wake effects. We found that a simple analytic approach derived from one-dimensional momentum theory, based on combining the wind speed distribution with the turbine thrust and power curves, explains 98% of the variance in simulated wake-induced losses. It follows that wake effects will be sensitive to the choice of turbine model, and therefore that location-dependent selection of turbine models may help to mitigate wake-induced losses.
@@
In assessing wind resources, either for early-stage wind farm planning or to derive inputs into power or energy system models, our results imply that it is not sufficient to assume spatially uniform wake-induced losses. The results also serve to highlight the natural advantages of northern Europe for offshore wind; this region has high ambient resources with relatively low wake-induced losses. In contrast, a number of emerging offshore wind markets are likely to be impacted by these enhanced (relative to mid-latitudes) wake-induced losses, including Brazil, Colombia, Australia and India.
@@
This study also explored the impact of wind direction distribution on the spatial variation in wind power potential. A narrower spread of wind directions presents greater opportunity for the optimisation of array layout designs. The narrowest wind direction distributions tend to be found in the tropics where the wind climate is characterised by the trade winds. Although we found that the narrower wind speed distributions found in these regions leads to stronger wake effects, the narrower wind direction distribution presents greater opportunity for mitigation of these losses via array layout optimisation. Although layout optimisation can only recover a fraction of the potential power lost to wakes, this result reveals the greater importance attached to layout optimisation in the tropics, compared with the mid-latitudes.
@@
Finally, we assessed future changes in wind power potential and wake-induced losses, due to climate change. Based on the SSP585 scenario for the years 2081–2100, compared with a ‘historical’ period of 1995–2014, much of the globe shows a small but significant decrease in ambient wind resource. This includes Europe and China, the two regions with the greatest current offshore wind capacity. The main exceptions to this are the polar regions and in particular the Arctic, where resources are projected to increase. As for the current climate, we find a correlation between changes in ambient wind resource, and percentage power lost to wakes. This means that regions where wind resources decrease due to climate change are also likely to suffer from greater percentage wake losses, exacerbating the climate effect. Therefore, in assessing climate impacts on wind resources, it is not sufficient to assume that wake-induced losses will remain the same under a changing climate.
@@
The results of this study have the potential to inform the transfer of knowledge from regions such as Europe, where offshore wind is relatively mature, to emerging markets such as in South America and Southeast Asia. The results highlight the different challenges likely to be experienced in efficiently harnessing wind resource in different regions, as well as the significance of layout optimisation, and potential future climate impacts.
@@
There are a number of avenues for further work. Firstly, this study considered only a single wind farm configuration, and in particular a single turbine model. As discussed above, this has an impact on the simulated wake-induced losses, and warrants further investigation. Secondly, the layout optimisation in this work has considered only a fixed number of turbines, and optimised the total power generated by the farm. Alternatively, optimising the levelised cost of energy with respect to a variable number of turbines might lead to mitigation of wake impacts, for example by reducing the installed capacity density in regions more strongly affected by wakes. Finally, we have focused here on intra-farm wake effects. Recent literature has highlighted the significance of inter-farm wake effects, which would add a further spatial (and directional) component to maps of wake-induced losses, as well as additional uncertainty with respect to unknown future build-out, which would need to be considered alongside uncertain climate change.
@@ Acknowledgements
Appendix A Linear regression model for optimisation potential
As described in section 2.3, to estimate the optimisation potential for our 100-turbine farm, we instead calculate the optimisation potential for a smaller 25-turbine farm, and train a model to predict the optimisation potential of a 100-turbine farm subject to the same wind rose.
The model is trained on a set of 288 wind roses. For each, we perform layout optimisation using TopFarm, for both the 25-turbine and 100-turbine wind farm cases. Figure A1 compares the optimisation potentials for each of these cases, showing a relationship which is close to linear. The 100-turbine farm tends to show slightly greater optimisation potential than the smaller 25-turbine farm. This is consistent with the larger farm experiencing stronger wake effects overall, and therefore having slightly greater scope for mitigation of wake-induced losses via optimisation.
To test the performance of a linear regression model to predict 100-turbine optimisation potential from 25-turbine optimisation potential, we divide the 288 optimisation cases into equal training and test datasets. The training dataset is used to train the model. Figure A2 compares the predicted and actual 100-turbine optimisation potentials for the test data. The root mean squared error is 0.4 percentage points. The accuracy of this model is sufficient to assess the overall global trends in optimisation potential.
Finally, the model is retrained using all 288 wind rose cases. The 25-turbine optimisation is performed for 18,491 grid cells with global coverage as described in section 2.3. The trained linear regression model is then used to predict the 100-turbine optimisation potential for all 18,491 grid cells.

### v2

@@ 4 Conclusions
This study has investigated global spatial variation in (i) ‘ambient’ offshore wind power potential, and (ii) intra-farm wake losses for a ‘standard’ offshore wind farm with a design representative of near-future installations. The spatial distribution of ambient wind power potential follows well-established global wind climatology, with resources greatest in the mid-latitudes, and lowest in the tropics. There is a strong correlation between ambient resources and percentage wake-induced losses; regions with greater ambient resource tend to lose a smaller fraction of their ambient power to wake effects. This trend exacerbates the spatial variation in ambient wind power density.
@@
However, there remains a broad distribution of wake-induced losses even amongst locations with similar ambient potentials. The tropics are particularly badly impacted by wake effects, with the mid-latitudes relatively less impacted. That is, for a tropical and mid-latitude location with the same ambient wind potential, the mid-latitude location is likely to suffer from smaller wake-induced losses than the tropical location. This can be explained by the specific wind speed distributions found in each region. The tropics tend to exhibit narrower wind speed distributions than mid-latitude regions, and thus a greater proportion of the wind speed distribution falls between the cut-in and rated speeds of the wind turbine, where power outputs are most sensitive to wake effects. A simple analytic approach derived from one-dimensional momentum theory, based on combining the wind speed distribution with the turbine thrust and power curves, explains 96% of the variance in simulated wake-induced losses. It follows that wake effects will be sensitive to the choice of turbine model, and therefore that location-dependent selection of turbine models may help to mitigate wake-induced losses.
@@
In assessing wind resources, either for early-stage wind farm planning or to derive inputs into power or energy system models, our results imply that it is not sufficient to assume spatially uniform wake-induced losses. The results also serve to highlight the natural advantages of northern Europe for offshore wind; this region has high ambient potential with relatively low wake-induced losses. In contrast, a number of emerging offshore wind markets are likely to be impacted by these enhanced (relative to mid-latitudes) wake-induced losses, including Brazil, Colombia, Australia and India.
@@
This study also explored the impact of wind direction distribution on the spatial variation in wind power potential. A narrower spread of wind directions presents greater opportunity for the optimisation of array layout designs. The narrowest wind direction distributions tend to be found in the tropics where the wind climate is characterised by the trade winds. Although the narrower wind speed distributions found in these regions lead to stronger wake effects, the narrower wind direction distribution presents greater opportunity for mitigation of these losses via array layout optimisation. Although layout optimisation can only recover a fraction of the potential power lost to wakes, this result reveals the greater importance attached to layout optimisation in the tropics, compared with the mid-latitudes.
@@
Finally, this study assessed future changes in wind power potential and wake-induced losses, due to climate change. Based on the SSP585 scenario for the years 2081–2100, compared with a ‘historical’ period of 1995–2014, much of the globe shows a small but significant decrease in ambient wind potential. This includes Europe and China, the two regions with the greatest current offshore wind capacity. The main exceptions to this are the polar regions and in particular the Arctic, where wind potential is projected to increase. As for the current climate, there is a correlation between changes in ambient wind potential, and percentage power lost to wakes. This means that regions where wind resources decrease due to climate change are also likely to suffer from greater percentage wake losses, exacerbating the climate change effect. Therefore, in assessing climate change impacts on wind potential, it is not sufficient to assume that wake-induced losses will remain the same under a changing climate.
@@
The results of this study have the potential to inform the transfer of knowledge from regions such as Europe, where offshore wind is relatively mature, to emerging markets such as in South America and Southeast Asia. The results highlight the different challenges likely to be experienced in efficiently harnessing wind resources in different regions, as well as the significance of layout optimisation, and potential climate change impacts.
@@
There are a number of avenues for further work. Firstly, this study considered only a single wind farm configuration, and in particular a single turbine model. As discussed above, this has an impact on the simulated wake-induced losses, and warrants further investigation. In particular, by repeating parts of this study for a variety of current or future turbine designs, opportunities to adapt turbine designs to particular wind climates could be explored. Secondly, the layout optimisation in this work has considered only a fixed number of turbines, and optimised the total power generated by the farm. Alternatively, optimising the levelised cost of energy (LCOE) with respect to a variable number of turbines might lead to further mitigation of wake impacts, for example by reducing the installed capacity density in regions more strongly affected by wakes. An LCOE-based optimisation would require a well-validated, globally applicable cost model, which may be a barrier to a global study, but region-specific studies may be feasible. Further cost considerations could also include the impact of increased turbulence due to wake effects on turbine maintenance costs and operating lifetime. Finally, the focus of this study is on intra-farm wake effects. Recent literature has highlighted the significance of inter-farm wake effects, which would add a further spatial (and directional) component to maps of wake-induced losses, as well as additional uncertainty with respect to unknown future build-out. A study of inter-farm wake effects and their uncertainty would require the generation of (an ensemble of) plausible wind farm build-out trajectories. The resulting uncertainties would also need to be considered alongside uncertain climate change.